OpenStudy (hba):

Sin6θ - sin 4θ + sin2θ = ?

OpenStudy (anonymous):

2 sin 4x .cos2x - sin 4x

OpenStudy (hba):

How ?

OpenStudy (anonymous):

sin a + sin b = 2 sin (a+b)/2 . cos (a-b)/2

OpenStudy (hba):

^ Ok

OpenStudy (hba):

Got this formula whats next

OpenStudy (anonymous):

nothing xD

OpenStudy (anonymous):

2 sin 4x .cos2x - sin 4x sin 4x ( 2cos 2x -1)

OpenStudy (anonymous):

u need to find general solution?

OpenStudy (hba):

I dont get it really :(

OpenStudy (anonymous):

whaat?? o.O

OpenStudy (hba):

Please Explain it to Me :(

OpenStudy (anonymous):

did u get this 2 sin 4x .cos2x - sin 4x

OpenStudy (hba):

No Thats Exactly what i didnt get, where did this come from ?

OpenStudy (anonymous):

lol using this sin a + sin b = 2 sin (a+b)/2 . cos (a-b)/2 sin 6x + sin 2x

OpenStudy (anonymous):

3sin (2x) is the simplified form.

OpenStudy (anonymous):

\[\sin(6x)-\sin(4x)+\sin(2x) =\sin(6x)+\sin(-4x)+\sin(2x)\] \[=2\sin(\frac{ 6x-4x }{ 2 })\cos(\frac{ 6x-4x }{ 2 })+\sin(2x)\] \[=2\sin(x)\cos(x)+\sin(2x)\] \[=2\sin(2x)+\sin(2x)\]\[=3\sin (2x)\]